# EVCap: Retrieval-Augmented Image Captioning with External Visual–Name Memory for Open-World Comprehension

### 27 Nov 2023 v1

@@ Abstract
Large language models (LLMs)-based image captioning has the capability of describing objects not explicitly observed in training data; yet novel objects occur frequently, necessitating the requirement of sustaining up-to-date object knowledge for open-world comprehension. Instead of relying on large amounts of data and scaling up network parameters, we introduce a highly effective retrieval-augmented image captioning method that prompts LLMs with object names retrieved from External Visual–name memory (EVCap). We build ever-changing object knowledge memory using objects’ visuals and names, enabling us to (i) update the memory at a minimal cost and (ii) effortlessly augment LLMs with retrieved object names utilizing a lightweight and fast-to-train model. Our model, which was trained only on the COCO dataset, can be adapted to out-domain data without additional fine-tuning or retraining. Our comprehensive experiments conducted on various benchmarks and synthetic commonsense-violating data demonstrate that EVCap, comprising solely 3.97M trainable parameters, exhibits superior performance compared to other methods of equivalent model size scale. Notably, it achieves competitive performance against specialist SOTAs with an enormous number of parameters. Our code is available at https://jiaxuan-li.github.io/EVCap.
@@ 1 Introduction
Advanced image captioning based on large language models (LLMs) [3, 9, 25, 8] has focused on the approach using big-scale models trained on ever-increasingly large-scale datasets, which is no longer viable. This is because the computational cost to train the models increases exponentially and, more importantly, updating training data is almost impossible that keep pace with the growth of novel objects in our daily lives. Sustaining ever-changing object knowledge with a reasonable cost is a pressing concern in LLMs-based models to truly unlock open-world comprehension.
@@
Figure 1: Overall comparison of our EVCap and SOTAs. (Upper) Comparison of the number of trainable parameters, CIDEr score on COCO and NoCaps datasets. The size of each circle reflects the log number of trainable parameters. EVCap (3.97M) has less trainable parameters than others while achieving comparable results with SOTAs at scale. (Lower) Generated captions by SmallCap, BLIP-2, and our EVCap for a commonsense-violating image from the WHOOPS dataset. $\times$ and $\checkmark$ indicate incorrect and correct predictions, respectively. Incorrect objects in captions are highlighted in red, while correct ones are in blue. SmallCap and BLIP-2 give incorrect predictions for “ice skates” and “wood floor”, respectively, while our EVCap utilizes an external visual–name memory to enhance attention to objects within the image, leading to superior performance for image captioning.
@@
Retrieval-augmented image captioning [35, 20] is emerging as an alternative since it considerably reduces training costs in both time and data while producing encouraging results. Nonetheless, with their huge datastore, it is obvious that LLMs would imitate the given texts, limiting their ability to describe open-world objects properly. For instance, SmallCap [35] considers the words “skateboard” and “wooden floor” to be a pair regardless of visual appearances containing a commonsense-violating pair of “ice skates” and “wood floor” (Fig. 1, lower). Additionally, prompting the LLMs given a lot of retrieved texts becomes cumbersome, requiring more trainable parameters. Fig. 1 (upper) shows that the CIDEr scores obtained by a lightweight SmallCap [35] with 43M trainable parameters are far away from those obtained by a heavy REVEAL [20] with 2.1B trainable parameters. Beyond that, due to the frequent occurrence of new objects, access to their sample texts is not always feasible, making the memory utilized in [35, 20] difficult to grow. We thus aim to streamline the external memory used in previous work [35, 20] by storing a sufficiently small amount of object information. And, of course, not only does the model not stereotype the example sentences, but the number of trainable parameters would be reduced drastically as a result of the causation (Fig. 1).
@@
We follow [40, 13] to construct a key-value memory where the key is represented by object’s features, and the value corresponds to object’s name. Unlike [40, 13], which rely on object definition as the key, our method leverages the visual appearance of the object as the key because of the abundance of object images readily available on the internet. We propose an external visual–name memory tailored for ease of expansion and cost-effectiveness in upholding up-to-date object information. We present a highly effective retrieval-augmented LLMs-based image captioning method, called EVCap, that prompts frozen LLMs with object names retrieved from our proposed memory for open-world comprehension. EVCap contains a frozen image encoder ViT [14] and Q-Former [25] with trainable image query tokens for object retrieval, an attentive fusion module, a trainable linear layer for mapping between vision and language latent spaces, and a frozen LLM decoder [10] for generating captions. Specifically, the attentive fusion module feeds retrieved object names and visual features into a customized frozen Q-Former using trainable object name query tokens to implicitly reduce the presence of superfluous object names. As a result, EVCap amounts to only 3.97M trainable parameters. Once trained, the model can be adapted to new domains and large-scale data without further fine-tuning or retraining. Our contributions are as follows:
@@
  • We present a remarkably lightweight and highly efficacious retrieval-augmented image captioning EVCap with 3.97M trainable parameters.
@@ 2 Related Work
Novel object captioning is a branch of image captioning that describes images containing objects that were not seen during training. Non-LLMs-based methods explore more objects by learning from unpaired image-sentence sources (DCC [19], NOC [39]) or relied on novel object detectors to recognize novel concepts (NBT [28], OSCAR [26] and VinVL [45]). LLMs-based methods such as ViECap [15] leverage the pre-trained CLIP [32] to obtain object entities. Nevertheless, the cut-off in training time of the pre-trained object detector or CLIP prevents it from detecting novel objects that arise quickly in reality. Unlike earlier work, we can readily update our recognition of novel concepts by adding them to external memory, ensuring that we keep any new objects from the past and even the future.
Retrieval-augmented image captioning is a recently popular approach that augments the captioning model with retrieved information for better open-world understanding. AoANet [16] uses a memory bank of image-sentence pairs and target words. SmallCap [35] employs image-to-text retrieval to obtain sampled captions from a captions datastore. RA-CM3 [44] retrieves documents from an external memory of a mixture of text and image via a dense multimodal retriever. EXTRA [34] and Re-ViLM [43] exploit the similarity of the input image and vision candidates to retrieve captions. Unlike previous methods, our external memory contains visual–name pairs to avoid redundant information in the external captions/documents. In addition, we use an attentive fusion module to mitigate the effects of irrelevant retrieved object names on caption generation.
@@ 3.1 Idea of EVCap
We aim to build a retrieval-augmented LLMs-based image captioning model with a sufficiently small yet informative external memory. It involves two challenges: (1) constructing an expandable external memory, and (2) building an effective LLMs-based model using retrieved object names.
@@
We base our method on a frozen pre-trained vision model and LLM with several trainable layers (Fig. 2), giving in a model that is cheap to train. To guide the LLM, we adopt a recently popular approach called prompting as in [29, 25, 11, 35, 46]. We begin by matching the learned visual features from the input image with image embeddings stored in memory, retrieving object names. We also introduce an attentive fusion module designed to implicitly remove irrelevant retrieved names. Finally, following the attentive fusion, we combine the learned visual features and object name features to form a prompt for the LLM to generate a caption, thus addressing challenge (2).
@@ 3.3 Object names retrieval
$\{\mathbf{k}_{j}^{\mathrm{best}},v_{j}^{\mathrm{best}}\}=\arg\max_{\mathbf{k}^{i}}\operatorname{SIM}\left(\mathbf{q}_{j},\mathbf{k}^{i}\right),\\ \vspace*{-0.75\baselineskip}$
@@ 3.4 Attentive fusion
The retrieved object names $\{v_{l}\}^{\mathrm{K}}_{l=1}$ are concatenated together into a sequence $\mathcal{S}$, each separated by a delimiter: $\mathcal{S}=\{v_{1},\texttt{[SEP]},v_{2},\texttt{[SEP]},\cdots,\texttt{[SEP]},v_{\mathrm{K}}\}$. The sequence $\mathcal{S}$ and visual features $\mathcal{Q}$ are fed into a customized Q-Former $\mathcal{F(\cdot)}$, which is constructed from the frozen pre-trained Q-Former as we used in vision encoder $\mathcal{E}$. Nonetheless, in order to enable object names to get attention from visual features, we switch the image embedding port and the text instruction port (see [25] for architecture detail). Like in the image encoding process in Sec. 3.3, we make the object name query tokens $\mathbf{T}_{\mathrm{obj}}$ learnable during training to assist in learning object name features related to the caption. The size of $\mathbf{T}_{\mathrm{obj}}$ is $P\times 768$, where $P$ indicates the number of object name query tokens. We get the object name features $\mathcal{V}=\mathcal{F}(\mathcal{S},\mathcal{Q},\mathbf{T}_{\mathrm{obj}})$.
@@ 4.1 Training setup
Training dataset. For all experiments, we exclusively train EVCap using the training set of COCO dataset [27], consisting of 82k images and 5 captions per images. The entire training process takes about 3 hours on 4 A6000 GPUs, using mixed precisions (more details in the supplementary).
@@ 5.1 Results on in-/out-domain benchmarks
Figure 3: Examples of captions generated by our EVCap and three SOTA methods on COCO test set, NoCaps validation set, and Flickr30k test set. GT refers to the Ground Truth captions. Incorrect objects in captions are highlighted in red, while correct ones are in blue. Our EVCap correctly generates captions across different datasets, showing performance comparable to BLIP-2. In contrast, SmallCap and MiniGPT4 sometimes either miss objects or include incorrect ones in their generated captions.
@@
Qualitative results. Fig. 3 presents a comparison of captions generated by our EVCap and three SOTA models across three benchmarks. The captions of SmallCap are generated by its publicly accessible demo [1]. We generate captions of MiniGPT4 and BLIP-2 using their respective pre-trained models. As a lightweight and retrieval-augmented captioning method, SmallCap struggles to produce accurate captions for given images, primarily because it relies on retrieved captions laden with extraneous information. MiniGPT4, though aligned with the primary content of images, sometimes misses certain objects like “trees” and “headphones”. This oversight stems from its focus on the main objects in images, without integrating additional cues for other objects provided by the retrieved object names. In contrast, the captions generated by our EVCap are comparable to those of BLIP-2.
@@ 5.2 Results on commonsense-violating data
Qualitative results. Fig. 4 illustrates the captions generated by EVCap, EVCap (w/WHOOPS), and three SOTAs for two images from the WHOOPS dataset. The first image challenges common sense as it unusually pairs “Einstein” with “racing car”. While all SOTAs simply refer to “Einstein” as “a man”, our EVCap and EVCap (w/WHOOPS) correctly identify him. The second image shows another unusual example. Similar to other methods except for BLIP-2, EVCap can not recognize “blue cartoon character” as “Pikachu”, while EVCap (w/WHOOPS) successfully predicts it because of the updated memory. In these two images, SmallCap and MiniGPT4 tend to generate captions with hallucinatory objects, a result of commonsense-violating contents present in the images.
@@ 5.3 Detailed analysis
Table 3: Ablation study on components prior to the LLM decoder in EVCap. The result of “+ Attentive fusion” demonstrates the substantial impact of the external visual–name memory.
@@ C S C S C S
Ablation study. We assess the contribution of each component prior to the LLM decoder in EVCap by incrementally integrating the image query tokens and the attentive fusion module into our baseline model. The baseline model comprises a ViT+Q-Former, a linear layer, and a LLM decoder. The quantitative results are shown in Tab. 3. When employing only the baseline model (Baseline), CIDEr scores drop notably by 5.7, 10.5, and 7.6 on COCO, NoCaps, and Flickr30k, respectively. The inclusion of trainable image query tokens (Baseline+) brings a marginal improvement on NoCaps and Flickr30k. However, the performance is significantly enhanced with the addition of attentive fusion (along with the introduction of external memory), indicating the pivotal role of the external visual–name memory in the overall effectiveness of EVCap. This is further corroborated by the qualitative results in Fig. 5, where captions from Baseline and Baseline+ inaccurately include objects like “couch” and “bed”, and Baseline+ overlooks “hand”.
@@
Figure 7: CIDEr scores after training EVCap with the number of retrieved object names $\mathrm{K}$ from 1 to 20. The results indicate that the performance is relatively optimal when $\mathrm{K}$ is set to be 10.
Impact of the number of retrieved object names. We investigate how the number of retrieved object names $\mathrm{K}$ (Sec. 3.3) affect EVCap in Fig. 7. We train the model with $\mathrm{K}$ from 1 to 20 and evaluate the performance under CIDEr on all three benchmarks. From the results, we can find that the model works worst on the out-domain dataset (NoCaps) when only one retrieved object name is used. As we gradually add more object names, performance fluctuates but improves. This pattern aligns with our intuition that when using one object name, the model will make errors or miss some objects in generated captions due to the incorrect object name. When we add more object names, we increase the error tolerance, and the attentive module in EVCap automatically pays more attention to image-related object names, thus improving results. Furthermore, we observe that setting $\mathrm{K}$ to 10 yields relatively optimal overall performance, validating the choice of $\mathrm{K}=10$ in EVCap.
@@
Limitations. First, EVCap cannot retrieve all objects that appear in the given image, leading to incomplete image descriptions as the second example in Fig. 4. We will investigate integrating object detection with image captioning to enhance completeness. Second, our focus on object representation restricts consideration of other crucial captioning elements, affecting overall performance. Similar to all models trained on the COCO dataset, our EVCap inherits its captioning style, which limits its ability to generate varied styles. This limitation is reflected in our relatively modest performance improvements in Tab. 2, compared to MiniGPT4. We will overcome this limitation by exploring methodologies that encourage style diversity in the future.

### 07 Apr 2024 v2

@@ Abstract
Large language models (LLMs)-based image captioning has the capability of describing objects not explicitly observed in training data; yet novel objects occur frequently, necessitating the requirement of sustaining up-to-date object knowledge for open-world comprehension. Instead of relying on large amounts of data and/or scaling up network parameters, we introduce a highly effective retrieval-augmented image captioning method that prompts LLMs with object names retrieved from External Visual–name memory (EVCap). We build ever-changing object knowledge memory using objects’ visuals and names, enabling us to (i) update the memory at a minimal cost and (ii) effortlessly augment LLMs with retrieved object names by utilizing a lightweight and fast-to-train model. Our model, which was trained only on the COCO dataset, can adapt to out-of-domain without requiring additional fine-tuning or re-training. Our experiments conducted on benchmarks and synthetic commonsense-violating data show that EVCap, with only 3.97M trainable parameters, exhibits superior performance compared to other methods based on frozen pre-trained LLMs. Its performance is also competitive to specialist SOTAs that require extensive training. *Equal contributions. Code is available at https://jiaxuan-li.github.io/EVCap.
@@ 1 Introduction
Advanced image captioning based on large language models (LLMs) [3, 9, 25, 8] has focused on the approach using big-scale models trained on ever-increasingly large-scale datasets, which is no longer viable. This is because the computational cost to train the models increases exponentially and, more importantly, updating training data is almost impossible to keep pace with the growth of novel objects in our daily lives. Sustaining ever-changing object knowledge with a reasonable cost is a pressing concern in LLMs-based models to truly unlock open-world comprehension.
@@
Figure 1: Overall comparison of our EVCap and SOTAs. (Upper) Generated captions by SmallCap, BLIP-2, and our EVCap for a commonsense-violating image from the WHOOPS dataset. $\times$ and $\checkmark$ indicate incorrect and correct predictions, respectively. Incorrect objects in captions are highlighted in red, while correct ones are in blue. SmallCap and BLIP-2 give incorrect predictions for “ice skates" and “wood floor", respectively, while our EVCap utilizes an external visual–name memory to enhance attention to objects within the image, leading to superior performance for image captioning. (Lower) Comparison of the number of trainable parameters, CIDEr score on COCO and NoCaps datasets. The size of each circle reflects the log number of trainable parameters. EVCap (3.97M) has less trainable parameters than others while achieving comparable results with SOTAs at scale.
@@
Retrieval-augmented image captioning [35, 20] is emerging as an alternative since it considerably reduces training costs in both time and data while producing encouraging results. Nonetheless, with their huge datastore, it is obvious that LLMs would imitate the given texts, limiting their ability to describe open-world objects properly. For instance, SmallCap [35] considers the words “skateboard" and “wooden floor" to be a pair regardless of visual appearances containing a commonsense-violating pair of “ice skates" and “wood floor" (Fig. 1, upper). Additionally, prompting the LLMs given a lot of retrieved texts becomes cumbersome, requiring more trainable parameters. Fig. 1 (lower) shows that the CIDEr scores obtained by a lightweight SmallCap [35] with 43M trainable parameters are far away from those obtained by a heavy REVEAL [20] with 2.1B trainable parameters. Beyond that, due to the frequent occurrence of new objects, access to their sample texts is not always feasible, making the memory utilized in [35, 20] difficult to grow. We thus aim to streamline the external memory used in previous work [35, 20] by storing a sufficiently small amount of object information. And, of course, not only does the model not stereotype the example sentences, but the number of trainable parameters would be reduced drastically as a result of the causation (Fig. 1).
@@
We follow [40, 13] to construct a key-value memory where the key is represented by object’s features, and the value corresponds to object’s name. Unlike [40, 13], which rely on object definition as the key, our method leverages the visual appearance of the object as the key because of the abundance of object images readily available on the internet. We propose an external visual–name memory tailored for ease of expansion and cost-effectiveness in upholding up-to-date object information. We present a highly effective retrieval-augmented LLMs-based image captioning method, called EVCap, that prompts frozen LLMs with object names retrieved from our proposed memory for open-world comprehension. EVCap contains a frozen image encoder ViT [14] and Q-Former [25] with trainable image query tokens for object retrieval, an attentive fusion module, a trainable linear layer for mapping between vision and language latent spaces, and a frozen LLM decoder [10] for generating captions. Specifically, the attentive fusion module feeds retrieved object names and visual features into a customized frozen Q-Former using trainable object name query tokens to implicitly reduce the presence of superfluous object names. As a result, EVCap amounts to only 3.97M trainable parameters. Once trained, the model can be adapted to new domains and large-scale data without further fine-tuning or re-training. Our contributions are as follows:
@@
  • We present a highly efficacious retrieval-augmented image captioning EVCap with 3.97M trainable parameters.
@@ 2 Related Work
Novel object captioning is a branch of image captioning that describes images containing objects that were not seen during training. Non-LLMs-based methods explore more objects by learning from unpaired image-sentence sources (DCC [19], NOC [39]) or rely on novel object detectors to recognize novel concepts (NBT [28], OSCAR [26] and VinVL [45]). LLMs-based methods such as ViECap [15] leverage the pre-trained CLIP [32] to obtain object entities. Nevertheless, the cut-off in training time of the pre-trained object detector or CLIP prevents it from detecting novel objects that arise quickly in reality. Unlike earlier work, we can readily update our recognition of novel concepts by adding them to external memory, ensuring that we keep any new objects from the past and even the future.
Retrieval-augmented image captioning is a recently popular approach that augments the captioning model with retrieved information for better open-world understanding. AoANet [16] uses a memory bank of image-sentence pairs and target words. SmallCap [35] employs image-to-text retrieval to obtain sampled captions from a captions datastore. RA-CM3 [44] retrieves documents from an external memory of a mixture of text and image via a dense multimodal retriever. EXTRA [34] and Re-ViLM [43] exploit the similarity of the input image and vision candidates to retrieve captions. Different from the previous methods, our external memory contains visual–name pairs to avoid redundant information in the external captions/documents. In addition, we use an attentive fusion module to mitigate the effects of irrelevant retrieved object names on caption generation.
@@ 3.1 Idea of EVCap
We aim to build a retrieval-augmented LLMs-based image captioning model with a sufficiently small yet informative external memory. It involves two challenges: (1) constructing an expandable external memory and (2) building an effective LLMs-based model using retrieved object names.
@@
We base our method on the frozen pre-trained vision model and LLM with several trainable layers (Fig. 2), giving in a model that is cheap to train. To guide the LLM, we adopt a recently popular approach called prompting as in [29, 25, 11, 35, 46]. We begin by matching the learned visual features from the input image with image embeddings stored in memory, retrieving object names. We also introduce an attentive fusion module designed to implicitly remove irrelevant retrieved names. Finally, following the attentive fusion, we combine the learned visual features and object name features to form a prompt for the LLM to generate a caption, thus addressing challenge (2).
@@ 3.3 Object names retrieval
$\{\mathbf{k}_{j}^{\mathrm{best}},v_{j}^{\mathrm{best}}\}=\arg\max_{\mathbf{k}^{i}}\operatorname{SIM}\left(\mathbf{q}_{j},\mathbf{k}^{i}\right),\$
@@ 3.4 Attentive fusion
The retrieved object names $\{v_{l}\}^{\mathrm{K}}_{l=1}$ are concatenated together into a sequence $\mathcal{S}$, each separated by a delimiter: $\mathcal{S}=\{v_{1},\texttt{[SEP]},v_{2},\texttt{[SEP]},\cdots,\texttt{[SEP]},v_{\mathrm{K}}\}$. The sequence $\mathcal{S}$ and visual features $\mathcal{Q}$ are fed into a customized Q-Former $\mathcal{F(\cdot)}$, which is constructed from the frozen pre-trained Q-Former as we used in vision encoder $\mathcal{E}$. Nonetheless, in order to enable object names to get attention from visual features, we switch the image embedding port and the text instruction port (see the supplement for architecture detail). Like in the image encoding process in Sec. 3.3, we make the object name query tokens $\mathbf{T}_{\mathrm{obj}}$ learnable during training to assist in learning object name features related to the caption. The size of $\mathbf{T}_{\mathrm{obj}}$ is $P\times 768$, where $P$ indicates the number of object name query tokens. We get the object name features $\mathcal{V}=\mathcal{F}(\mathcal{S},\mathcal{Q},\mathbf{T}_{\mathrm{obj}})$.
@@ 4.1 Training setup
Training dataset. For all experiments, we exclusively train EVCap using the training set of COCO dataset [27], consisting of 82k images and 5 captions per images. The entire training process takes about 3 hours on 4 A6000 GPUs, using mixed precisions (more details in the supplement).
@@ 5.1 Results on in-/out-domain benchmarks
Figure 3: Examples of captions generated by our EVCap and three SOTA methods on COCO test set, NoCaps validation set, and Flickr30k test set. GT refers to the Ground Truth captions. Incorrect objects in captions are highlighted in red, while correct ones are in blue. Our EVCap correctly generates captions across different datasets, showing performance comparable to BLIP-2.
@@
Qualitative results. Fig. 3 presents a comparison of captions generated by our EVCap and three SOTA models across three benchmarks. The captions of SmallCap are generated by its publicly accessible demo [1]. We generate captions of MiniGPT4 and BLIP-2 using their respective pre-trained models. As a lightweight and retrieval-augmented captioning method, SmallCap struggles to produce accurate captions for given images, primarily because it relies on retrieved captions laden with extraneous information. MiniGPT4, though aligned with the primary content of images, sometimes misses certain objects like “trees" and “headphones". This oversight stems from its focus on the main objects in images, without integrating additional cues for other objects provided by the retrieved object names. In contrast, the captions generated by our EVCap are comparable to those of BLIP-2.
@@ 5.2 Results on commonsense-violating data
Qualitative results. Fig. 4 illustrates the captions generated by EVCap, EVCap (w/ WHOOPS), and three SOTAs for one image from the WHOOPS dataset. Similar to other methods except for BLIP-2, EVCap cannot recognize “blue cartoon character" as “Pikachu", while EVCap (w/WHOOPS) successfully predicts it because of the updated memory. SmallCap and MiniGPT4 tend to generate captions with hallucinatory objects, a result of commonsense-violating contents present in the images.
@@ 5.3 Detailed analysis
Table 3: Ablation study on components prior to the LLM decoder in EVCap. The result of “+ Attentive fusion" demonstrates the substantial impact of the external visual–name memory.
@@ C S C S C S
Ablation study. We assess the contribution of each component prior to the LLM decoder in EVCap by incrementally integrating the image query tokens and the attentive fusion module into our baseline model. The baseline model comprises a ViT+Q-Former, a linear layer, and a LLM decoder. The quantitative results are shown in Tab. 3. When employing only the baseline model (Baseline), CIDEr scores drop notably by 5.7, 10.5, and 7.6 on COCO, NoCaps, and Flickr30k, respectively. The inclusion of trainable image query tokens (Baseline+) brings a marginal improvement on NoCaps and Flickr30k. However, the performance is significantly enhanced with the addition of attentive fusion (along with the introduction of external memory), indicating the pivotal role of the external visual–name memory in the overall effectiveness of EVCap. This is further corroborated by the qualitative results in Fig. 5, where captions from Baseline and Baseline+ inaccurately include objects like “couch" and “bed", and Baseline+ overlooks “hand".
@@
Figure 7: CIDEr scores after training EVCap with the number of retrieved object names $\mathrm{K}$ from 0 to 20. The results indicate that the performance is relatively optimal when $\mathrm{K}$ is set to be 10.
Impact of the number of retrieved object names. We investigate how the number of retrieved object names $\mathrm{K}$ (Sec. 3.3) affect EVCap in Fig. 7. We train the model with $\mathrm{K}$ from 0 to 20 and evaluate the performance under CIDEr on all three benchmarks. From the results, we can find that the model works worst on the out-domain dataset (NoCaps) with zero object names. It confirms that the performance boost from Baseline+ to EVCap (Tab. 3) is primarily attributed to the retrieval-augmented mechanism, but not the customized Q-Former itself. With more object names, performance fluctuates but improves. Furthermore, we observe that setting $\mathrm{K}$ to 10 yields relatively optimal overall performance, validating the choice of $\mathrm{K}=10$ in EVCap.
@@
Limitations.
First, EVCap cannot retrieve all objects that appear in the given image due to the memory coverage limits, leading to incomplete image descriptions (Fig. 4). We will investigate integrating object detection with image captioning to enhance completeness. Second, our focus on object representation restricts consideration of other crucial captioning elements, affecting overall performance. Similar to all models trained with COCO dataset, EVCap has limitations in generating varied styles, which is reflected in our relatively modest performance improvements in Tab. 2, compared to MiniGPT4. We will overcome it by exploring methodologies that encourage style diversity in the future.